OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3).

OpenStudy (anonymous):

its 11 just substitute 3 in the place of x. 3^2 + 2 = 9+2 =11

OpenStudy (anonymous):

\[g(\Box)=\Box^2+2\] put \(3\) in the box

OpenStudy (anonymous):

Thank you(:

OpenStudy (anonymous):

ur welcome